OpenStudy (anonymous):

Simplify the expression. 7 sinh(ln t)

OpenStudy (mathmale):

I'd suggest you look up the definition of the hyperbolic sine as a function of the variable x. It involves the exponential function. Once you have that definition, replace each occurrence of x with ln t and simplify. Note that e^ln p = p (as an example). Finally, multiply your result by 7 and you'll be done.

OpenStudy (anonymous):

I approached the problem a different way. I used the chain rule. f(t)=sinh(t) g(t)=ln(t) f'(t)=cosh(t) g'(t)=1/t So that would give me --> 7cosh(lnt)*1/x because of the fact that the chain rule states that f'g(x)*g'(x). I do not know if this is correct though.

OpenStudy (anonymous):

oh sorry i accidentally put x as the denominator, but it's supposed to be t

OpenStudy (perl):

the directions did not say to take the derivative , though

OpenStudy (perl):